# Characterizing the MrDeepFakes Sexual Deepfake Marketplace

### 14 Oct 2024 v1

@@ Abstract
The prevalence of sexual deepfake material has exploded over the past several years. Attackers create and utilize deepfakes for many reasons: to seek sexual gratification, to harass and humiliate targets, or to exert power over an intimate partner. In tandem with this growth, several markets have emerged to support the buying and selling of sexual deepfake material. In this paper, we systematically characterize the most prominent and mainstream marketplace, MrDeepFakes. We analyze the marketplace economics, the targets of created media, and user discussions of how to create deepfakes, which we use to understand the current state-of-the-art in deepfake creation. Our work uncovers little enforcement of posted rules (e.g., limiting targeting to well-established celebrities), previously undocumented attacker motivations, and unexplored attacker tactics for acquiring resources to create sexual deepfakes.
@@ 1 Introduction
The prevalence of deepfakes—media created or manipulated with artificial intelligence—has exploded over the last several years [29]. Most deepfakes are nonconsensual sexual media: recent reports estimate that 98% of online deepfakes are pornographic [20]. Beyond sexual gratification, attackers have also created and disseminated sexual deepfakes to humiliate or harass a target [^1] Drawing upon prior work on online harassment, we use “target” over “victim” to avoid disempowering those facing abuse. [1] or to exert control over an intimate partner [32]. These sexual deepfakes threaten the emotional, psychological, and physical well being of those targeted [19].
@@
This growth has been supported by dedicated markets that have emerged for buying and selling sexual deepfake content. These digital marketplaces are often accessible on the open Internet and enable buyers to commission custom sexual deepfakes from technically savvy sellers. The most prominent and mainstream platform is MrDeepFakes, a site where visitors can view celebrity sexual deepfakes, buy and sell deepfakes, and engage in community discussion on how to create deepfakes. Prior work has examined the community’s migration away from Reddit, where it originated as r/deepfakes before being banned, to its new independent platform on mrdeepfakes.com [57]. However, no work has studied this ecosystem’s market dynamics, content produced, or technical capabilities since the explosion of deepfakes post-2021.
@@
In this paper, we systematically characterize the MrDeepFakes platform over the past five years. We document the rise of this online community and analyze the economics of the marketplace, focusing on the buyers, sellers, and targets of deepfakes posted to the site. Through a detailed analysis of forum discussions, we also examine the state-of-the-art technical components required to create deepfakes. We use mixed methods to analyze the data, combining comprehensive quantitative measurements of forum activity and video attributes with qualitative coding of forum threads.
@@
At the time of our study, MrDeepFakes hosted 43K sexual deepfake videos depicting 3.8K individuals; these videos have been watched more than 1.5 billion times. Analyzing the most common targets, who account for 95% of videos on the site, we find that 95.3% of individuals are women actors and musicians, making up 88% of posted videos. While MrDeepFakes claims to only serve media depicting “celebrity” targets, we find these rules unenforced and that hundreds of targeted individuals have little to no online or public presence. Similarly, despite purported rules banning “negative” depictions of targets, we find that 2.4% of uploaded videos include scenes of rape, abuse, or humiliation.
@@
We next examine 830 paid request threads on the forum-based marketplace. Using thematic analysis, we surface three underlying motives for buyers: nonconsensual sexual gratification, power and masculinity, and community contribution. We identify the centering of community and the creation of deepfakes as a perceived form of contributed art as a novel incentive that falls outside existing taxonomies of motives for online abuse [56, 9]. We identify 58 posts that contain prices that buyers and sellers were willing to pay or accept we find the average price of a deepfake video is \$87.50; prices ranged greatly and were as inexpensive as \$1 and as high as \$1,500. However, community guidance suggests that buyers should be wary of offers below \$50. Availability of data for the target (“facesets”), output quality, difficulty of angles, and turnaround time are the primary factors that contribute to cost.
@@
Finally, we examine $1229$ posts across 25 deepfake creation guide and tutorial threads to study the current state-of-the-art. We document the three most crucial components that pose challenges for deepfake creators: access to facial image data, software proficiency, and powerful hardware. We find that the MrDeepFakes community relies on a host of resources, from open source software to free-tier cloud GPU access (e.g., Google Colab notebooks). We highlight that the community often references and utilizes recent academic research from well-established machine learning and computer vision venues.
@@
Ultimately, our work characterizes the sexual deepfake marketplace and documents the resources, challenges, and community-driven solutions that arise in the sexual deepfake creation process. We conclude with a discussion of our results, contextualizing implications of a new incentive for online abuse, considerations for the open source and academic communities, and suggestions for improving mitigations with legislation. We hope our work illuminates the shape of deepfake abuse and its underlying community and provides useful context for researchers, industry players, and policymakers to address this growing threat.
@@ 2 Background & Related Work
Our study is rooted in two bodies of related work: (1) interdisciplinary explorations of deepfake technology and the online communities that have formed around it, and (2) investigations of underground markets.
@@ 2.1 Deepfakes and Image-Based Sexual Abuse
Literature across computer science, economics, and women’s studies has sought to understand various aspects of deepfakes and other forms of image-based sexual abuse (IBSA), including online deepfake communities [57, 15, 62], societal perceptions of deepfake consumption and production [5, 36, 62], and motives for generating IBSA [9, 27, 25, 56, 54].
@@
Most relevant to our work, Timmerman et al. investigated the emergence of the MrDeepFakes forum in response to the 2018 deplatforming of r/deepfakes from Reddit [57]. Their work makes two primary contributions. First, they analyze how the community situates itself in society, such as its members’ perceptions of the ethics, legality, and social acceptability of deepfakes. They show that deplatforming events have sowed distrust of platforms and the media within the community, positioning MrDeepFakes as the only viable source of technical support for creating sexual deepfakes. Second, they characterize the targets of datasets shared within the community for use in deepfake creation, finding that they are primarily women celebrities from English-speaking countries. Although our work also studies MrDeepFakes, we focus on analyzing the economics of its deepfake marketplace, the explosion of deepfake content, and the deepfake creation process (e.g., requisite software, hardware, data, and expertise).
@@
Similarly, Widder et al. showed that core contributors and users of an unnamed, “safe for work” (SFW) open source deepfake tool believed that denying “valuable technical support” for pornographic usage was their best means of control for its ethical use [62]. In contrast, despite societal disapproval of creating and sharing deepfake nonconsensual intimate imagery (NCII) [5], another study of deepfake discussion on Reddit concluded that online subcommunities are eager to facilitate both creating and sharing deepfakes [15].
@@
Despite past findings that depict deepfake creation as challenging, Mehta et al. found that nonexperts can create deepfakes with relative ease; however, the poor realism of the participant-generated deepfakes indicates that creating photorealistic deepfakes remains difficult for novices [36]. Still, while consumers of deepfake NCII often know it is fake and “do not care,” the harms experienced by their targets—emotional and psychological trauma, reputational damage, and financial loss [12, 3, 41]—are independent of deepfake content’s achieved realism [6].
@@
Beyond deepfakes, extensive prior work has sought to understand different aspects of online abuse—the underlying motives of perpetrators, targets’ experiences of harm, and their risk factors—producing several frameworks and taxonomies [56, 59, 54, 35, 46, 18]. For deepfakes and other forms of IBSA, studies have identified motives including power, circumvention of consent, revenge, and financial gain [9, 56, 25, 27]. Using an economics framing, Kshetri introduced a cost-benefit model for abusive deepfake creation that accounts for investment, opportunity, and psychological costs and financial and psychological benefits [27]. We draw on these frameworks to interpret nonconsensual sexual deepfakes using a scoped threat model [56] with three stakeholders: the consumer, the creator, and the target. We define the target as one whose likeness is abused out of context; however, sex workers whose bodies nonconsensually appear in the videos are also harmed and is an area for future work.
@@ 2.2 Underground Forums and Marketplaces
The security community has previously studied online underground forums and markets, including the sale of stolen credentials and drugs [38, 55], bulletproof hosting services [40], and ransomware [16]. The profit-driven framing of attackers in such ecosystems has allowed researchers to model them as rational actors, using economic principles to identify potential disruption tactics. Tseng et al. apply similar methods to partner infidelity and surveillance forums, examining how members carry out intimate partner surveillance [58]. They observe that these communities openly discuss and develop new abuse strategies in public forums, concluding that improving moderation has mitigation potential [58].
Researchers have also identified vulnerabilities in the monetization infrastructure of cybercrime operations (e.g., sex trafficking, ransomware, spam) to leverage for the detection, attribution, or discouragement of such activity. These potential disruption channels include Bitcoin transactions [48], cryptocurrency exchanges [16], and policies of payment processors and online advertising providers [25, 17, 30]. This body of work demonstrates the promise of mitigations informed by measurements of underground forums and markets; we employ similar approaches to study the nonconsensual sexual deepfake ecosystem.
@@ 2.3 MrDeepFakes Platform
Figure 1: MrDeepFakes—The MrDeepFakes platform is composed of two spaces: a tube site for watching deepfake sexual content and a forum where buyers can solicit content.
MrDeepFakes is the largest celebrity sexual deepfake community [37, 34] and consists of two primary spaces: (1) a media sharing platform where members upload deepfake videos and images, and (2) a forum where they discuss deepfake-related topics (e.g., how to create high-quality deepfakes) and negotiate the exchange of deepfake-related commodities (e.g., videos and datasets) (Figure 1). The former is served as the landing page of MrDeepFakes as a “tube” site [^2] A pornographic video-sharing site (e.g., PornHub) where visitors can watch sexual deepfake videos. The forum is split into nine sections, including a marketplace and deepfake creation tool sections (Table 1).
3 Methodology
Our study is based on crawling the public MrDeepFakes website using headless Chrome and the Selenium WebDriver. We specifically collected public forum posts and replies, metadata attached to public videos, and public user profile information in the fall and winter of 2023. We created an account to access public site data, but we did not bypass any security controls, attempt to access any private data, download any video content, or interact with forum members. We never attempt to de-anonymize or investigate individual buyers or sellers. We describe the ethical implications of our data collection and analysis in greater detail in Section Ethics.
@@ 3 Methodology
Starting at PAGE=1, we paginated over /videos/[PAGE] until we hit an HTTP 404 Not Found error. This yielded URLS for $42986$ videos that we crawled to collect each video’s title, description, creator, measures of engagement, tagged individuals (deepfaked celebrities), and any provided Bitcoin address for donations to the uploader.
@@
We scraped each forum section landing page and collected URLs to each subsection and thread, which we scraped to collect the title and body along with associated responses, including message timestamp, author, and attachments. In total, we scraped 8,198 threads, comprising 43,350 forum posts. Table 1 shows a breakdown of each crawled forum section, and we provide the hierarchical section structure in Appendix A.
@@
Table 1: Forum Sections—Breakdown of activity (by thread count) in each section of the forum.
@@
Opening a MrDeepFakes account creates two public profiles by default: one on the tube site and one on the forum. On the forum, users are assigned monotonic user IDs, which we crawled by visiting /forums/members/[USER_ID] up to the highest known user ID (the forum links the profile of the newest user at any given time on its homepage). For each profile, we scraped the ID, banners (which denote special roles like moderator), profile posts, and bio. We collected data about 611K public profiles.
@@ 3.1 Limitations
Our analysis is limited to publicly available data, which does not capture private activity on the site, or communication that transitions to third-party platforms; we also exclude discourse in languages other than English. Therefore, what we report likely underestimates the degree of activity and does not capture more fringe and illicit behaviors occurring on MrDeepFakes, much less for the broader online deepfake community. Regarding data collection, the churn in forum, video, and user profile content meant that some data was not available during our scrape. We validated the completeness of our data by comparing it to the content on the website as rendered in our browser shortly after our data collection, and we manually verified the correctness of 10% samples of forum, video, and user data.

### 02 Jan 2025 v2

@@ Abstract
The prevalence of sexual deepfake material has exploded over the past several years. Attackers create and utilize deepfakes for many reasons: to seek sexual gratification, to harass and humiliate targets, or to exert power over an intimate partner. In tandem with part supporting this growth, several markets markets have emerged to support the buying and selling of sexual deepfake material. In this paper, we systematically characterize the most prominent and mainstream marketplace, MrDeepFakes. We analyze the marketplace economics, the targets of created media, and user discussions of how to create deepfakes, which we use to understand the current state-of-the-art in deepfake creation. Our work uncovers little enforcement of posted rules (e.g., limiting targeting to well-established celebrities), previously undocumented attacker motivations, and unexplored attacker tactics for acquiring resources to create sexual deepfakes.
@@ 1 Introduction
The prevalence of deepfakes—media created or manipulated with artificial intelligence—has exploded over the last several years [deepfakes-everywhere-bloomberg]. Most deepfakes are nonconsensual sexual media: recent reports estimate that 98% of online deepfakes are pornographicsexual in nature [homesecurityheroes2023state]. Beyond sexual gratification, attackers have also created and disseminated sexual deepfakes to humiliate or harass a targettargets [^1] Drawing upon prior work on online harassment, we use “target” over “victim” to avoid disempowering those facing abuse. [rana-ayyub-deepfake] or to exert control over an intimate partnerintimate partners [lucas2022deepfakes]. These sexual deepfakes threaten the emotional, psychological, and physical well being of those targetedwell-being of targets [hao21horrifying].
@@
This growth has been is supported by dedicated markets that have emerged for buying and selling sexual deepfake content. These digital marketplaces are often accessible on the open Internet and enable buyers to commission custom sexual deepfakes from technically savvy sellers. The most prominent and mainstream platform is MrDeepFakes, a site where visitors can view celebrity sexual deepfakes, buy and sell deepfakes, and engage in community discussion on how to create deepfakes. Prior work has examined the migration of this communitycommunity’s migration away from Reddit, where it originated as r/deepfakes before being banned, to its new independent platform on mrdeepfakes.com [timmerman2023studying]. However, no work has studied this ecosystem’s market dynamics, content producedhosted content, or technical capabilities since the explosion of deepfakes post 2021. post-2021.
@@
In this paper, we systematically characterize the MrDeepFakes platform over the past five years. We document the rise of this online the community and analyze the economics of the its marketplace, focusing on the buyers, sellers, and targets of deepfakes posted to the site. Through a detailed analysis of forum discussions, we also examine the state-of-the-art technical components required to create deepfakes. We use mixed methods to analyze the data, combining posted deepfakes. Combining comprehensive quantitative measurements of forum activity and video attributes with qualitative coding analysis of forum threads. , we answer the following three research questions:
  1. 1.
    How has sexual deepfake video creation and consumption changed over time?
  2. 2.
    What are the marketplace economics of buying and selling deepfake videos?
  3. 3.
    What are the requirements and challenges for creating deepfakes?
@@
At the time of our studyAs of November 2023, MrDeepFakes hosted 43K sexual deepfake videos depicting 3.8K individuals; the these videos have been watched more than 1.5 billion times. Analyzing the most common targets, who account for 95% of videos on the site, we find that 95.3% of individuals are women and actors and musiciansmake , making up 88% of posted videos. While MrDeepFakes claims to only serve media depicting “celebrity” targets, we find these rules unenforced and that hundreds of targeted individuals have little to no online or public presence. Similarly, despite purported rules banning “negative” depictions of targets, we find that 2.4% of uploaded videos include scenes of rape, abuse, or humiliation.
@@
We next examine 830 paid request threads on the forum-based marketplace. Using thematic analysis, we surface three underlying motives for buyers: nonconsensual sexual gratification, power and masculinity, and community contribution. We identify the centering of community and the creation of deepfakes perception of commissioning deepfakes as contributing art as a perceived form of contributed art as a novel incentive that falls outside existing taxonomies of motives for online abuse [thomas2021hatesok, okolie2023ibsa]. We
While most request threads do not specify a price, we identify 58 posts that contain prices that buyers and sellers were willing to pay or acceptand . We find the average price of a deepfake video is \$87.50with prices ranging ; prices ranged greatly and were as inexpensive as \$1 and as high as \$1,500. However, community guidance suggests that buyers should be wary of offers below \$50. Availability of data for the target (“facesets”), output quality, difficulty of angles, and turnaround time are the primary factors that contribute to costpricing.
@@
Finally, we examine $1229$ posts across 25 guide deepfake creation guides and tutorial threads for deepfake creation to study the current state-of-the-artin sexual deepfake creation. We document the three most crucial components that pose challenges for deepfake creators: access to facial image data, software proficiency, and powerful hardware. We find that the MrDeepFakes community relies on a host of resourcesranging , from open source software to free-tier cloud GPU access (e.g., Google Colab notebooks). We highlight that the community often references and utilizes recent academic research from well established well-established machine learning and computer vision venues . and bypasses protections that developers have attempted to build into their tools to prevent abuse.
@@
Ultimately, our work characterizes the sexual deepfake marketplace and documents the resources, challenges, and community-driven solutions that arise in the sexual deepfake creation process. We conclude with a discussion of our results, calling out contextualizing implications of a new incentive for online abuse, considerations for the open source and academic communities, and suggestions for improving mitigations with legislation . We hope our work helps to shine light on this community and issue and provides useful context to how legislation may potentially help. By illuminating the largest public community behind creating and sharing deepfakes, we hope to provide valuable context for researchers, industry playerspartners, and policymakers in addressing working to address this growing threat.
@@ 2 Background & Related Work
Our study is rooted in two bodies of related work: (1) interdisciplinary literature exploring explorations of deepfake technology and the online communities that have formed around it, and (2) prior investigations of underground markets.
@@ 2.1 Deepfakes and Image-Based Sexual Abuse
Literature across computer science, economics, and women’s studies has sought to understand various aspects of deepfakes and other forms of image-based sexual abuse (IBSA), including online deepfake communities [timmerman2023studying, gamage2022deepfakes, widder2022ethicalai], societal perceptions of deepfake consumption and production [brigham2024violation, mehta2023can, widder2022ethicalai], and motives for generating perpetrating IBSA [okolie2023ibsa, kshetri2023economics, bellingcat2023any, thomas2021hatesok, scheuerman2021framework].
@@
Most relevant directly related to our work, Timmerman et al. investigated the emergence of the MrDeepFakes forum in response to the 2018 deplatforming of r/deepfakes from Reddit [timmerman2023studying]. Their work makes two primary contributions. First, they analyze how the community situates itself in society, such as its members’ perceptions of the ethics, legality, and social acceptability of deepfakes. They show that deplatforming events have sowed distrust of mainstream platforms and the media within the community, positioning MrDeepFakes as the only viable source of technical support for creating sexual deepfakes. Second, they characterize the targets of datasets shared within the community for use in deepfake creation, finding that they are primarily women celebrities from English-speaking countries. Although our work also studies MrDeepFakes, we focus on analyzing the economics of its deepfake marketplace, the explosion of deepfake content, and the deepfake creation process (e.g., requisite software, hardware, data, and expertise).
@@
Similarly, Widder et al. showed found that core contributors and users of an unnamed, “safe for work” (SFW) open source deepfake tool believed that denying “valuable technical support” for pornographic usage usage in sexual contexts was their best means of control for its ethical use [widder2022ethicalai]. In contrast, despite societal disapproval of creating and sharing deepfake NCIInonconsensual intimate imagery (NCII) [brigham2024violation], another study of deepfake discussion on Reddit concluded that online subcommunities are eager to facilitate both creating and sharing deepfakes [gamage2022deepfakes].
@@
Despite past findings that depict deepfake creation as challenging, Mehta et al. found that nonexperts can create deepfakes with relative ease; however, the poor realism of the participant-generated deepfakes indicates that creating photorealistic deepfakes remains difficult for novices [mehta2023can]. Still, while consumers of deepfake NCII often know it is fake and “do not care,” the harms experienced by their targets—emotional and psychological trauma, reputational damage, and financial loss [dickson2020tiktok, ayyub2018silence, dhsincreasing]—are independent of deepfake content’s achieved realism [burkell2019nothing].
@@
Beyond deepfakes, extensive prior work has sought to understand different aspects of online abuse—the underlying motives of perpetrators, targets’ experiences of harm, and their risk factors—producing several frameworks and taxonomies [thomas2021hatesok, warford2022framework, scheuerman2021framework, marwick2021networked, phillips2019exclusionary, han2023hate]. For deepfakes deepfake NCII and other forms of IBSA, studies have identified motives including power, circumvention of consent, revenge, and financial gain [okolie2023ibsa, thomas2021hatesok, bellingcat2023any, kshetri2023economics]. Using an economics framing, Kshetri introduced a cost-benefit model for abusive deepfake creation that accounts for investment, opportunity, and psychological costs and financial and psychological benefits [kshetri2023economics]. We draw on these frameworks to interpret nonconsensual sexual deepfakes deepfake NCII using a scoped threat model [thomas2021hatesok] with three stakeholders: the consumer, the creator, and the target. We define the target as one whose likeness is abused out of context; however, sex workers, whose bodies nonconsensually appear in the videos, are also harmedand is an area , and this remains a direction for future work.
@@ 2.2 Underground Forums and Marketplaces
The security community has previously studied online underground forums and markets, including the sale of stolen credentials and drugs [motoyama11analysis, soska15measuring], bulletproof hosting services [noroozian19platforms], and ransomware [gray2022money]. The profit-driven framing of attackers in such ecosystems has allowed researchers to model them as rational actors, using economic principles to identify potential disruption tactics. Tseng et al. apply similar methods to partner infidelity and surveillance forums, examining how members carry out intimate partner surveillance [tseng20tools]. They observe that these communities openly discuss and develop new abuse strategies in public forums, concluding that improving moderation has mitigation potential [tseng20tools].
Researchers have also identified vulnerabilities in the monetization infrastructure of cybercrime operations (e.g., sex trafficking, ransomware, spam) to leverage for the detection, attribution, or discouraging disincentivization of such activity. These potential disruption channels include Bitcoin transactions [portnoff17backpage], cryptocurrency exchanges [gray2022money], and policies of payment processors and ad payment processing and online advertising providers [bellingcat2023any, han2022infra, levchenko2011trajectories]. This body of work demonstrates the promise of mitigations informed by measurements of underground forums and markets; we employ similar approaches to study the nonconsensual sexual deepfake deepfake NCII ecosystem.
2.3 Platform
Figure 1: MrDeepFakes—The MrDeepFakes platform is composed of two spaces: a tube site for watching deepfake sexual deepfake content and a forum where buyers can solicit content.
@@ 2.3 MrDeepFakes Platform
MrDeepFakes is the largest celebrity open sexual deepfake community [meikle2022deepfakes, maiberg22clicks] and consists of two primary spaces: (1) a media sharing platform where members upload deepfake videos and images; and deepfakes, and (2) a forum where they discuss for discussing deepfake-related topics (e.g., how to create high-quality deepfakes) and negotiate the exchange of deepfake-related commodities (e.g., videos and datasets) (Figure 1). The former is served as the landing page of MrDeepFakes as a “tube” site [^2] A pornographic video-sharing site (e.g., PornHub) where visitors can watch sexual deepfake videos. The forum is split into nine sections, including a marketplace and deepfake creation tool sections (Table 1).
3 MethodologyData Collection
Our study is based on crawling the public MrDeepFakes website using headless Chrome and the Selenium WebDriver. We specifically collected public forum posts and replies, metadata attached to public videos, and public user profile information in the fall and winter October and November of 2023. We created an account to access public site data, but we did not bypass any security controls, attempt to access any private data, download any video content, or interact with forum members. We never attempt to de-anonymize or investigate individual buyers or sellers. We describe the ethical implications of our data collection and analysis in greater detail in Section EthicsEthical Considerations.
@@ 3 MethodologyData Collection
To understand posted deepfake videos, starting
Deepfake Videos. Starting at PAGE=1, we paginated over /videos/[PAGE] until we hit an HTTP 404 Not Found error. This yielded URLS URLs for $42986$ videos that we crawled to collect each video’s title, description, creator, measures of engagement, tagged individuals (deepfaked celebrities), and any provided Bitcoin address for donations to the uploader.
@@
To understand ’s marketplace and community, we crawled the site’s forum.
Forum Posts.We scraped each forum section landing page and collected URLs to each subsection and thread, which we scraped to collect the title and body along with associated responses, including message timestamp, author, and attachments. In total, we scraped 8,198 threads, comprising 43,350 forum posts. Table 1 shows a breakdown of each crawled forum section, and we provide the hierarchical section structure in Appendix A.
@@
Table 1: Forum Sections—Breakdown —We present the breakdown of activity (by thread count ) in each section of the MrDeepFakes forum section.
@@
User Accounts.Opening a MrDeepFakes account creates two public profiles by default: one on the tube site and one on the forum. On the forum, users are assigned monotonic user IDs, which we crawled by visiting /forums/members/[USER_ID] up to the highest known user ID (the forum links the profile of the newest user at any given time on its homepage). For each profile, we scraped the ID, banners (which denote special roles like moderator), profile posts, and bio. We collected data about 611K public profiles.
@@ 3.1 Limitations
Our analysis is limited to publicly available data, which does not capture private activity on the site , or communication that transitions to or communication on third-party platforms; we also exclude discourse in languages other than Englishnon-English discourse. Therefore, what we report our work likely underestimates the degree of forum activity and does not capture more fringe and illicit behaviorsoccurring on , much less for the broader online deepfake community. Regarding data collection, the churn . Churn in forum, video, and user profile content meant that some data was not available during our scrapealso prevented some data from being collected scraped. We validated the completeness of our data by comparing it to the content on the website as rendered in our browser shortly after our data collection, and we manually verified the correctness of 10% samples of forum, video, and user data.